OpenStudy (anonymous):

1. Determine whether the following problems (a-c) show direct variation. If NOT, how could you change it to make it a direct variation? Remember, direct variation takes the form of y=kx. a. x 1 3 5 7 9 y 4 10 16 22 28 b. y=10x c. y=3x-4

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: direct variation between two variables \(x,\;y\), is modeled by the subsequent equation: \(y=kx\) for example

OpenStudy (anonymous):

i don't get it :c

OpenStudy (michele_laino):

namely, we can say y varies directly with x, if the subsequent equation holds: \(y=kx\), where \(k\) is any number, more precisely any real number

OpenStudy (mathmale):

" k " is known as the " proportionality constant." What are your learning resources? Do you have an algebra textbook? Do you have access to online learning materials for your algebra course? Please at least glance through some of this material to find "direct proportion" or "direct variation." Note that y = 7x represents direct proportion, whereas y = 7 x + 1 does not.

OpenStudy (anonymous):

@Michele_Laino is there an easy way or can you give an example

OpenStudy (michele_laino):

for example, equation c) is similar to this equation \(y=kx\) ?

OpenStudy (michele_laino):

for example a direct variation can b modeled by this formula: \(y=2x\)

OpenStudy (michele_laino):

can be*

OpenStudy (michele_laino):

whereas an equation like this: \(y=2x+1\), doesn't represent a direct variation, since it is present the constant value +1, at the right side

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

so, is option c) a direct variation?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

that's right! so, if we cancel the constant -4, at the right side, we get this formula: \(y=3x\) and \(y=3x\) is a direct variation?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

that's right! next, is option b) a direct variation?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

that's right, since such equation has this form: \(y=kx\), with \(k=10\)

OpenStudy (michele_laino):

now, for part a), we have to write the equation between x and y, using the data into the table above

OpenStudy (anonymous):

okay :D

OpenStudy (michele_laino):

in order to do that, I consider the first ordered pair \(x=1,\;y=4\), and I replace such values ito this formula: \(y=ax+b\), where \(a,\;b\) are two constant which have to be determined

OpenStudy (michele_laino):

so, after a substitution, I get this: \(4=a \cdot 1+b\) or: \(4=a+b\)

OpenStudy (anonymous):

so 4=a*1+b

OpenStudy (michele_laino):

yes! \(4=a+b\)

OpenStudy (michele_laino):

then I consider the second ordered pair: \(x=3,\;y=10\), and a fter a substitution, I get: \(10=a \cdot 3+b\), or: \(10=3a+b\)

OpenStudy (anonymous):

10=a*3+b

OpenStudy (michele_laino):

yes! so, we have to solve thys algebraic system: \[\left\{ \begin{gathered} 4 = a + b \hfill \\ 10 = 3a + b \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

oops.. this*

OpenStudy (anonymous):

hehe lol its okay

OpenStudy (michele_laino):

do you know how to solve such system?

OpenStudy (anonymous):

um i don't :c sorry

OpenStudy (michele_laino):

please wait a moment, I have to answer to my phone

OpenStudy (anonymous):

Okay :D

OpenStudy (michele_laino):

here I am

OpenStudy (anonymous):

Yay !! :)

OpenStudy (michele_laino):

we can solve such system, using this procedure: I subtract, side by side, the first equation from the second one, so I get: \[10 - 4 = 30 + b - a - b\]

OpenStudy (michele_laino):

oops.. I have made a typo: \[10 - 4 = 3a + b - a - b\]

OpenStudy (michele_laino):

then I simplify both sides, so I get: \[6 = 2a\]

OpenStudy (michele_laino):



OpenStudy (michele_laino):

next I divide both sides by 2, so I get: \[\frac{6}{2} = \frac{{2a}}{2}\] what is \(a\) ?